# ChartAB: A Benchmark for Chart Grounding & Dense Alignment

### v1

@@ 3.1 Dataset Taxonomy and Construction
Figure 2: Statistics of ChartAB. ChartAB includes 9,000 paired chart images curated for tasks below: (1) Paired charts for Data Grounding & Alignment differ in one to three data cells; (2) Paired charts for Attribute Grounding & Alignment differ in color, legend position, or text style; (3) Robustness task includes multiple pairs that share identical differences in data but differ in certain attributes.
@@ A.3 Dataset Construction
Input: Source dataset $\mathbf{D}_{\text{ChartX}}=\{(T_{i},S_{i})\}_{i=1}^{N}$ from ChartX (Xia et al., 2024), where $T_{i}$ is a CSV table and $S_{i}$ is the corresponding plotting script; Number of cells to modify $k\in\{1,2,3\}$; Scaling range $\alpha_{\min},\alpha_{\max}]$.
Output: Constructed dataset $\mathbf{D}^{(\mathrm{data})}_{\texttt{ChartAB}}=\{(x_{i},x^{\prime}_{i},y^{g}_{i},y^{a}_{i})\}_{i=1}^{M}$, where $x_{i},x^{\prime}_{i}$ are chart images, $y^{g}_{i}$ is the grounding label, and $y^{a}_{i}$ is the alignment label.
@@
Input: Source dataset $\mathbf{D}_{\text{ChartX}}=\{(T_{i},S_{i})\}_{i=1}^{N}$ from ChartX (Xia et al., 2024), where $S_{i}$ is the plotting script; Set of attribute types $\mathcal{A}=\{\text{color},\text{legend},\text{text style}\}$.
Output: Constructed dataset $\mathbf{D}_{\texttt{ChartAB}}^{\text{(attribute)}}=\{(x_{i},x^{\prime}_{i},at_{i},y^{g}_{i},y^{a}_{i})\}_{i=1}^{M}$, where $x_{i}$, $x^{\prime}_{i}$ are chart images, $at_{i}\in\mathcal{A}$: attribute type, $y^{g}_{i}$ is the grounding label, $y^{a}_{i}$ is the alignment label.
@@
 foreach attribute type $at\in\mathcal{A}$ do
@@
    if $at=\text{color}$ then
@@
    else if $at=\text{legend}$ then
@@
    else if $at=\text{text style}$ then
@@
         Create instance $(x,x^{\prime},at,y^{g},y^{a})$;
         Append $(x,x^{\prime},at,y^{g},y^{a})$ to $\mathbf{D}_{\texttt{ChartAB}}^{\text{(attribute)}}$;
@@
Input: Source dataset $\mathbf{D}_{\text{ChartX}}=\{(T_{i},S_{i})\}_{i=1}^{N}$; Data modification params: $k$, $\alpha_{\min},\alpha_{\max}]$; Visual variations per instance: $d=5$; Attribute types: $\mathcal{A}=\{\text{color},\text{legend},\text{text style}\}$.
Output: $\mathbf{D}_{\texttt{ChartAB}}^{\text{(robust)}}=\{\{(x^{(j)}_{i},x^{\prime(j)}_{i})\}_{j=1}^{d},y^{g}_{i},y^{a}_{i},at_{i}\}_{i=1}^{M}$ where $x^{(j)}_{i}$, $x^{{}^{\prime}(j)}_{i}$ are chart images for variation $j$, $at_{i}\in\mathcal{A}$: attribute type being varied, $y^{g}_{i}$ is the grounding label, $y^{a}_{i}$ is the alignment label.
foreach $at\in\mathcal{A}$ do
@@
         Sample variation $\Delta_{j}$ for attribute atat (color/legend/text style);
@@
         Append $\{\{(x^{(j)}_{i},x^{\prime(j)}_{i})\}_{j=1}^{d},y^{g}_{i},y^{a}_{i},at_{i}\}$ to $\mathbf{D}_{\texttt{ChartAB}}^{\text{(robust)}}$;
@@ A.5.1 Dense Alignment
where $\rho(\cdot,\cdot)\in[0,1]$ is a task-specific value matching function: exact match for categorical values (e.g., color hex codes, text styles), and $\rho(v,\hat{v})=1-\min(|v-\hat{v}|/|v|,1)$ for numerical values (e.g., data points, font sizes).
@@ A.5.2 Robustness
$r(x,x^{\prime})=\frac{1}{\sigma\left(\left\{s^{(\text{data})}_{\text{align}}(x^{(j)},x^{\prime(j)})\right\}_{i=1}^{d}\right)}$ (5)
@@ A.6.2 Ablations
1-stage Multi-chart 5.1 6.6 3.7 3.8 3.7 2.6 0.7 2.3 2.6
Stitched-chart 5.8 6.6 4.2 4.0 2.6 1.1 1.0 1.6 2.2
2-stage Ours 6.7 7.5 3.3 5.9 6.6 2.5 2.2 1.4 2.1
@@
The ablation experiments analyzed Phi-3.5 model’s performance on data alignment task. As shown in table. 3, the single-stage approach fared poorly compared to out two-stage approach reaffirming the two-stage approach. Multi-image inference showed the weakest performance. Despite increasing training efforts towards improved VLM training, the models still face issues in reasoning ability on fine-grained tasks. Stitched-charts approach showed better results than multi-image, however they too underperformed vis-a-vis our two-stage approach. The comparatively stronger image self-attention capabilities seem to augment multi-image by utilzing the stitched connection. However the better prevailing capabilities of two-stage approach capture the gain of grounding generation. The VLM’s multi-modal understanding though improving still suffers from finer-level nuances missed by information loss in image-encoding and cross-attention mechanisms.
@@ A.7 Additional Finding & Insights
Figure 12: (a) Legend alignment of legend positions. Phi-3.5 performs the worst while GPT-4o is best. Related discussion in Finding 1&2. (b) Text-style alignment (size, weight, font). Worst: QWEN-2.5-VL, Best: GPT-4o. Discussion in Finding 1&4.

### v2

@@ 3.1 Dataset Taxonomy and Construction
Figure 2: Statistics of ChartAB. ChartAB includes 9,000+ instances curated for tasks below: (1) Paired charts for Data Grounding & Alignment differ in one to three data cells; (2) Paired charts for Attribute Grounding & Alignment differ in color, legend position, or text style; (3) Robustness task includes multiple pairs that share identical differences in data but differ in certain attributes.
@@ A.3 Dataset Construction
Input: Source dataset $\mathbf{D}_{\text{ChartX}}=\{(T_{i},S_{i})\}_{i=1}^{N}$ from ChartX (Xia et al., 2024), where $T_{i}$ is a CSV table and $S_{i}$ is the corresponding plotting script, NN is number of instances; Number of cells to modify $k\in\{1,2,3\}$; Scaling range $\alpha_{\min},\alpha_{\max}]$.
Output: Constructed dataset $\mathbf{D}^{(\mathrm{data})}_{\texttt{ChartAB}}=\{(x_{i},x^{\prime}_{i},y^{g}_{i},y^{a}_{i})\}_{i=1}^{M}$, where $x_{i},x^{\prime}_{i}$ are chart images, $y^{g}_{i}$ is the grounding label, and $y^{a}_{i}$ is the alignment label, $M$ is number of instances.
@@
Input: Source dataset $\mathbf{D}_{\text{ChartX}}=\{S_{i}\}_{i=1}^{N}$ from ChartX (Xia et al., 2024), where $S_{i}$ is the plotting script, NN is number of instances; Set of attribute types $B=\{\text{color},\text{legend},\text{text style}\}$.
Output: Constructed dataset $\mathbf{D}_{\texttt{ChartAB}}^{\text{(attribute)}}=\{(x_{i},x^{\prime}_{i},b_{i},y^{g}_{i},y^{a}_{i})\}_{i=1}^{M}$, where $x_{i}$, $x^{\prime}_{i}$ are chart images, $b_{i}\in B$ is the attribute type, $y^{g}_{i}$ is the grounding label, $y^{a}_{i}$ is the alignment label, $M$ is number of instances.
@@
 foreach attribute type $b\in B$ do
@@
    if $b=\text{color}$ then
@@
    else if $b=\text{legend}$ then
@@
    else if $b=\text{text style}$ then
@@
         Create instance $(x,x^{\prime},b,y^{g},y^{a})$;
         Append $(x,x^{\prime},b,y^{g},y^{a})$ to $\mathbf{D}_{\texttt{ChartAB}}^{\text{(attribute)}}$;
@@
Input: Source dataset $\mathbf{D}_{\text{ChartX}}=\{(T_{i},S_{i})\}_{i=1}^{N}$, where $T_{i}$ is a CSV table and $S_{i}$ is the corresponding plotting script, NN is number of instances; Number of cells to modify $k\in\{1,2,3\}$; Scaling range $\alpha_{\min},\alpha_{\max}]$; Visual variations per instance $d=5$; Set of attribute types: $B=\{\text{color},\text{legend},\text{text style}\}$.
Output: $\mathbf{D}_{\texttt{ChartAB}}^{\text{(robust)}}=\{\{(x^{(j)}_{i},x^{\prime(j)}_{i})\}_{j=1}^{d},y^{g}_{i},y^{a}_{i},at_{i}\}_{i=1}^{M}$ where $x^{(j)}_{i}$, $x^{{}^{\prime}(j)}_{i}$ are chart images for variation $j$, $b_{i}\in B$ is the attribute type being varied, $y^{g}_{i}$ is the grounding label, $y^{a}_{i}$ is the alignment label.
foreach $b\in B$ do
@@
         Sample variation $\Delta_{j}$ for attribute $b$ (color/legend/text style);
@@
         Append $\{\{(x^{(j)},x^{\prime(j)})\}_{j=1}^{d},y^{g},y^{a},at\}$ to $\mathbf{D}_{\texttt{ChartAB}}^{\text{(robust)}}$;
@@ A.5.1 Dense Alignment
where $\rho(\cdot,\cdot)\in[0,1]$ denotes a task-specific value matching function: it performs exact matching for categorical attributes (e.g., text weight/font), $\rho(v,\hat{v})=1-\|v-\hat{v}\|_{2}$ for color attributes (with $\|v-\hat{v}\|_{2}$ denoting the normalized RGB distance), and $\rho(v,\hat{v})=1-\min(|v-\hat{v}|/|v|,1)$ for numerical attributes (e.g., data values or text size).
@@ A.5.2 Robustness
$r(x,x^{\prime})=\frac{1}{1+\sigma\left(\left\{s^{(\text{data})}_{\text{align}}(x^{(j)},x^{\prime(j)})\right\}_{i=1}^{d}\right)}$ (5)
@@ A.6.2 Ablations
1-stage Multi-chart 2.6 4.5 1.9 2.9 3.0 1.1 0.1 0.9 0.8
Stitched-chart 2.1 2.2 0.8 1.9 0.9 0.5 0.1 0.1 0.4
2-stage Ours 4.7 7.0 1.7 5.4 5.9 1.0 0.1 0.4 0.7
@@
The ablation experiments evaluated the Phi-3.5 model’s performance on the data alignment task. As shown in Table 3, the single-stage approach underperformed compared to our proposed two-stage method, reaffirming the effectiveness of intermediate grounding for reasoning in the alignment task, helping to focus more precisely on localized relationships between visual and textual elements. In contrast, the single-stage approach struggles to capture these fine-grained correspondences due to information loss during joint encoding and limited cross-attention resolution. Despite continued progress in multi-modal training, current VLMs still face challenges in detailed reasoning, and our results highlight how decomposing complex tasks like our multi-chart dense alignment into modular stages can substantially mitigate these limitations.
@@ A.7 Additional Finding & Insights
Figure 12: (a) Legend alignment of legend positions. Phi-3.5 performs the worst while GPT-4o is best. Related discussion in Finding 1&2. (b) Text-style alignment (size, weight, font). Worst: InternVl-2.5-8B, Best: GPT-4o. Discussion in Finding 1&4.